# Comparison of chiral limit studies in curvature mass versus on-shell renormalized quark-meson model using ChPT

### v1

@@ I Introduction
The FRG QM Model chiral limit study, has estimated the impact of quark loop vacuum fluctuations on the extent of crossover, second and first order transition regions of the Columbia plot, after switching off the thermal and quantum fluctuations of mesons in the Wetterich equation wetrich guided FRG flow of the scale-dependent effective action. They have emphasized Resch that by switching off the meson loops, they are directly working in the extended mean field approximation (e-MFA) within the FRG which is same as the alternative method of including quark one-loop vacuum term in the QM Model where the vacuum divergences are regularized by the standard dimensional regularization technique and curvature masses of the mesons are used to fix the model parameters as in the Ref. schafwag12 . The above direct treatment of the fermionic vacuum term in the QM model, has a long history and a large number of such QMVT model studies Herbst ; schafwag12 ; vac ; Gatto ; Dima ; Anna ; lars ; guptiw ; chatmoh1 ; vkkr12 ; TranAnd ; chatmoh2 ; vkkt13 ; Weise1 ; Weyrich ; kovacs ; Weise2 ; zacchi1 ; zacchi2 ; Rai ; Weise3 ; Gyozo , have already quantified the strong influence of the fermionic vacuum fluctuations on the QCD phase diagram and the strength of the chiral phase transition occurring at the physical point for experimental $\pi,K,\eta\text{ and }\eta^{\prime}$ masses. It is important to generate the Columbia plot in the $m_{\pi}-m_{K}$ and quark mass $m_{ud}-m_{s}$ planes of the 2+1 flavor QMVT model by using the $\mathcal{O}(\frac{1}{f^{2}})$ accurate infrared regularized U(3) ChPT Herpay:05 ; gasser ; borasoyI ; borasoyII ; Beisert ; Becher scaling relations for the ($m_{\pi},m_{K}$) dependence of $f_{\pi},f_{K}\text{ and }M_{\eta}$ such that the QMVT model parameter fixing, away from the physical point becomes free from any ambiguity or heuristic adjustment when the light strange chiral symmetry breaking source strengths are reduced towards the chiral limit.
@@
The double derivatives of the effective potential at its minimum, define the curvature masses of the mesons. Hence the quark one-loop vacuum corrections to meson self energies are included at the zero momentum in the curvature masses as the effective potential is the generator of the n-point functions of the theory at vanishing external momenta. Noticing the above, another approach of treating the quark one-loop vacuum fluctuations consistently, similar to the work in Ref. laine , was recently developed in Refs. Adhiand1 ; Adhiand2 ; Adhiand3 ; asmuAnd ; RaiTiw ; raiti2023 for the two flavor and in Refs. vkkr22 ; skrvkt24 for the 2+1 flavor of the QM model where one finds the on-shell renormalized QM (RQM) model parameters by matching the counterterms in the modified minimal subtraction $\overline{MS}$ scheme with the corresponding counterterms in the on-shell scheme when the running couplings and mass parameter are put into the relation of the physical (pole) masses of the $\pi,K,\eta,\eta^{\prime}\text{ and }\sigma$ mesons. Very recently, QM model extension with diqaurks, has been studied in Ref. Ander25 ; Gholami where the above method of on-shell renormalization of model parameters is applied. We investigated the chiral limit and generated the $m_{\pi}-m_{K}$ ( $m_{ud}-m_{s}$) planes of the Columbia plot, in our very recent 2+1 flavor RQM model study vkt25 where two input sets of the $\mathcal{O}(\frac{1}{f^{2}})$ accurate results from the infrared regularized U(3) ChPT Herpay:05 ; gasser ; borasoyI ; borasoyII ; Beisert ; Becher and the standard large $N_{c}$ U(3) ChPT gasser ; LeutI ; KaisI ; herrNPB ; herrPLB ; Escribano , were used for obtaining the ($m_{\pi},m_{K}$) dependence of the $f_{\pi},f_{K}\text{ and }M_{\eta}$ when one moves away from the physical point. Though the $\eta^{\prime}$ meson is treated differently in the two ChPT approaches, only marginal differences were found in the extent of first, second order and crossover transition regions of the Columbia plot when the RQM model results from two input sets were compared in Ref. vkt25 . Hence, use of only the infrared regularized U(3) ChPT input, will serve the purpose for computing the QMVT model Columbia plot as proposed in the above paragraph.
@@
Comparing the QMVT model Columbia plots in the present work calculated for the $m_{\sigma}=400\text{ and }530$ MeV with the corresponding Columbia plots in the RQM Model and e-MFA:QM model FRG study (for $m_{\sigma}=530$ MeV), we will estimate how the first, second or crossover chiral transition regions in the $m_{\pi}-m_{K}$($m_{ud}-m_{s}$) and the $\mu-m_{K}$($\mu-m_{s}$) planes, get qualitatively and quantitatively modified by the different methods of treating the quark one-loop vacuum fluctuations in the QM model. Since e-MFA:QM and the QMVT model both use curvature masses of mesons to fix the model parameters after implementing the dimensional regularization of the one-loop vacuum divergence, the differences in the first, second order and crossover transition regions of the Columbia plots, will give us the estimate of the differences which arise due to the fact that the $f_{\pi},f_{K},M_{\eta}$ and parameters, change towards the chiral limit in the QMVT model (guided by the ChPT given $(m_{\pi}-m_{K})$ dependence), whereas the $f_{\pi},~{}M_{\eta}$ and parameters in the e-MFA:QM-FRG study are kept fixed to their physical point value and the initial effective action is successively adjusted to the higher scales $\Lambda$ as one approaches the chiral limit. The physical point scalar $\sigma$ vacuum mass of $m_{\sigma}=530$ MeV decreases significantly in the e-MFA:QM-FRG study while it does not change in the QMVT model when one moves away from the physical point to reach the chiral limit. In complete contrast to the above two scenarios, the RQM model $\sigma$ meson vacuum curvature mass $m_{\sigma,c}$ has minimum value at the physical point which increases towards the chiral limit whereas the pole mass $m_{\sigma}=530$ MeV remains the same. The computed $m_{\sigma}$ temperature variations will be compared across the models when $m_{\pi}-m_{K}$ are reduced in going to the chiral limit.
@@ II The Model formulations
The Lagrangian Rischke:00 ; Schaefer:09 ; TiPQM3F of QM model is
@@
The three flavor of quark fields $\psi$ (color $N_{c}$-plet Dirac spinor) are coupled by the Yukawa coupling $g$ to the nine scalar(pseudo-scalar) meson $\xi$ fields $\sigma_{a}(\pi_{a}$) of $3\times 3$ complex matrix ${\cal{M}}=T_{a}\xi_{a}=T_{a}(\sigma_{a}+i\pi_{a})$$T_{a}=\frac{\lambda_{a}}{2}$ where $\lambda_{a}$ ($a=0,1..8$) are Gell-Mann matrices with $\lambda_{0}=\sqrt{\frac{2}{3}}{\mathbb{I}}_{3\times 3}$. The field $\xi$ picks up the non-zero vacuum expectation value $\overline{\xi}$ in the 00 and $8$ directions. The condensates $\bar{\sigma_{0}}$ and $\bar{\sigma_{8}}$ break the $SU_{L}(3)\times SU_{R}(3)$ chiral symmetry spontaneously while the external fields $H=T_{a}h_{a}$ with $h_{0}$, $h_{8}\neq 0$ break it explicitly. The change from the singlet octet $(0,8)$ to the non-strange strange basis $(x,y)$ gives $x(h_{x})=\sqrt{\frac{2}{3}}\bar{\sigma}_{0}(h_{0})+\frac{1}{\sqrt{3}}\bar{\sigma}_{8}(h_{8})$ and $y(h_{y})=\frac{1}{\sqrt{3}}\bar{\sigma}_{0}(h_{0})-\sqrt{\frac{2}{3}}\bar{\sigma}_{8}(h_{8})$. Considering mesons at mean field level with the thermal and quantum fluctuations of the quarks/anti-quarks, the grand potential Schaefer:09 ; TiPQM3F , is the sum of the vacuum effective potential $U(x,y)$ and the quark/anti-quark contribution $\Omega_{q\bar{q}}$ at finite temperature $T$ and quark chemical potential $\mu_{f}(f=u,d,s)$.
@@ II.1 QMVT Model
The derivation of the QMVT model effective potential, is briefly described below. The curvature masses of the scalar and pseudo-scalar mesons, are used to fix the parameters in the QMVT model when the quark one-loop vacuum fluctuation in the Eq. (6) is included in its effective potential under the the extended mean field approximation (e-MFA). The vacuum divergence of the Eq. (6) is regularized under the minimal subtraction scheme using the dimensional regularization as it was done in the Refs. vac ; guptiw ; vkkr12 for the two flavor and in the Refs. schafwag12 ; chatmoh1 ; vkkt13 for the three flavor. After the dimensional regularization of the Eq. (6) near three dimensions, $d=3-2\epsilon$, one gets :
${\hskip-2.84544pt}\Omega_{q\bar{q}}^{\rm vac}=\sum_{f=u,d,s}\frac{N_{c}\ m_{f}^{4}}{16\pi^{2}}\left[\frac{1}{\epsilon}-\frac{\{-3+2\gamma_{E}+4\ln(\frac{m_{f}}{2\sqrt{\pi}\Lambda})\}}{2}\right],$ (9)
@@
The $x$ and $y$ dependent curvature masses of the mesons determine the six unknown parameters $m^{2}$, $\lambda_{1},\lambda_{2}$, $h_{x},h_{y}$ and $c$ in the meson potential U($x,y$). The details of the parameter fixing can be found in the Ref. vkkt13 . The parameter $\lambda_{2}$ is determined as $\lambda_{2}=\lambda_{2s}+n+\lambda_{2+}+\lambda_{2\Lambda}$ where $\lambda_{2s}$ is the same old $\lambda_{2}$ parameter of the QM/PQM model in the Ref. Rischke:00 ; Schaefer:09 ; TiPQM3F $n=\frac{N_{c}g^{4}}{32\pi^{2}}$, $\lambda_{2+}=\frac{n{f_{\pi}}^{2}}{f_{K}\left(f_{K}-f_{\pi}\right)}\ln\{\frac{2f_{K}-f_{\pi}}{f_{\pi}}\}$ and $\lambda_{2\Lambda}=4n\ln\{\frac{g\left(2f_{K}-f_{\pi}\right)}{2\Lambda}\}$. The renormalization scale $\Lambda$ dependent part $\lambda_{2\Lambda}$ in the above is generated by the logarithmic $\Lambda$ dependence in the term $\Omega_{q\bar{q}}^{\rm vac}$. After substituting this value of the $\lambda_{2}$ in the expression of U($x,y$) and writing explicitly all the terms under the summation in the expression of $\Omega_{q\bar{q}}^{\rm vac}$, when one does the rearrangement of the terms, one finds that the scale dependence of all the terms in $\Omega_{q\bar{q}}^{\rm vac}$ gets completely canceled by the logarithmic $\Lambda$ dependence of the $\lambda_{2}$ contained in $\lambda_{2\Lambda}$. Finally, the scale independent expression of the vacuum effective potential is obtained as :
@@
The $\lambda_{2\text{v}}=\lambda_{2s}+\lambda_{2+}$. Note that the quark one-loop vacuum corrections, in the QMVT model parameter fixing scheme, modify the parameters $m^{2}$, $\lambda_{1}$ and $\lambda_{2}$ and the unaffected parameters $h_{x}$, $h_{y}$ and $c$ remain the same as in the QM model. The QMVT model thermodynamic grand potential is written as
@@ II.2 RQM model
$-2\frac{c_{0}}{g^{3}_{0}}\Delta_{x}^{2}\ \Delta_{y}+4\frac{\lambda_{10}}{g^{4}_{0}}\Delta_{x}^{2}\ \Delta_{y}^{2}+2\frac{(2\lambda_{10}+\lambda_{20})}{g^{4}_{0}}\Delta_{x}^{4}$
@@
The terms $\mathcal{C}(m^{2}_{\pi})$ and $\mathcal{C}^{\prime}(m^{2}_{\pi})$ and the derivations for the renormalized parameters $m^{2}_{0}=(m^{2}+m^{2}_{\text{\tiny{FIN}}})$$h_{x0}=(h_{x}+h_{x\text{\tiny{FIN}}})$$h_{y0}=(h_{y}+h_{y\text{\tiny{FIN}}})$$\lambda_{10}=(\lambda_{1}+\lambda_{1\text{\tiny{FIN}}})$$\lambda_{20}=(\lambda_{2}+\lambda_{2\text{\tiny{FIN}}})$ and $c_{0}=(c+c_{\text{\tiny{FINTOT}}})$ are given in detail in Refs. vkkr22 ; skrvkt24 . The $m^{2}_{\text{\tiny{FIN}}}$, $h_{x\text{\tiny{FIN}}}$, $h_{y\text{\tiny{FIN}}}$, $\lambda_{1\text{\tiny{FIN}}}$, $\lambda_{2\text{\tiny{FIN}}}$ and $c_{\text{\tiny{FINTOT}}}$ are the finite on-shell corrections in the parameters at the scale $\Lambda_{0}$. The experimental values of the pseudo-scalar meson masses $m_{\pi}$, $m_{K}$, $m_{\eta}$, $m_{\eta^{\prime}}$, the scalar $\sigma$ mass $m_{\sigma}$ and the $f_{\pi}$, $f_{K}$ as input determine the tree level QM model quartic couplings $\lambda_{1}$, $\lambda_{2}$, mass parameter $m^{2}$$h_{x}$, $h_{y}$ and the coefficient $c$ of the t’Hooft determinant term for the $U_{A}(1)$ axial anomaly Rischke:00 ; Schaefer:09 .
@@
+$+\frac{\left\{2(\lambda_{1}+\lambda_{1\text{\tiny{FIN}}})+(\lambda_{2}+\lambda_{2\text{\tiny{FIN}}})\right\}x^{4}}{8}+\left(\lambda_{1}+\lambda_{1\text{\tiny{FIN}}}+\lambda_{2}\right.$
$\left.+\lambda_{2\text{\tiny{FIN}}}\right)\frac{y^{4}}{4}+\frac{N_{c}g^{4}(x^{4}+2y^{4})}{8(4\pi)^{2}}\left[\frac{3}{2}-\mathcal{C}(m^{2}_{\pi})-m^{2}_{\pi}\mathcal{C}^{\prime}(m^{2}_{\pi})\right]$
@@
The search of the grand potential minima $\frac{\partial\Omega_{\rm RQM}}{\partial x}=\frac{\partial\Omega_{\rm RQM}}{\partial y}=0$ for the Eq. (19) gives the $T$ and $\mu$ dependence of the $x$ and $y$. The curvature masses of mesons are different from their pole masses in the RQM model due to the on-shell renormalization of the parameters BubaCar ; fix1 . The derivation of the expressions of curvature masses of the mesons are presented in the section (IIA) of the Ref.(vkt25, ). Using the Eq. (II.2) to evaluate the equations of motion $\frac{\partial\Omega_{vac}^{\rm RQM}}{\partial x}=0=\frac{\partial\Omega_{vac}^{\rm RQM}}{\partial y}$, one gets the renormalized explicit chiral symmetry breaking strengths $h_{x0}$ and $h_{y0}$ as the following.

### v2

@@ I Introduction
The FRG QM Model chiral limit study, has estimated the impact of quark loop vacuum fluctuations on the extent of crossover, second and first order transition regions of the Columbia plot, after switching off the thermal and quantum fluctuations of mesons in the Wetterich equation wetrich guided FRG flow of the scale-dependent effective action. They have emphasized Resch that by switching off the meson loops, they are directly working in the extended mean field approximation (e-MFA) within the FRG which is same as the alternative method of including quark one-loop vacuum term in the QM Model where the vacuum divergences are regularized by the standard dimensional regularization technique and curvature masses of the mesons are used to fix the model parameters as in the Ref. schafwag12 . The above direct treatment of the fermionic vacuum term in the QM model, has a long history and a large number of such QMVT model studies Herbst ; schafwag12 ; vac ; Gatto ; Dima ; Anna ; lars ; guptiw ; chatmoh1 ; vkkr12 ; TranAnd ; chatmoh2 ; vkkt13 ; Weise1 ; Weyrich ; kovacs ; Weise2 ; zacchi1 ; zacchi2 ; Rai ; Weise3 ; Gyozo , have already quantified the strong influence of the fermionic vacuum fluctuations on the QCD phase diagram and the strength of the chiral phase transition occurring at the physical point for experimental $m_{\pi},m_{K},m_{\eta}\text{ and }m_{\eta^{\prime}}$. It is important to generate the Columbia plot in the $m_{\pi}-m_{K}$ and quark mass $m_{ud}-m_{s}$ planes of the 2+1 flavor QMVT model by using the $\mathcal{O}(\frac{1}{f^{2}})$ accurate infrared regularized U(3) ChPT Herpay:05 ; gasser ; borasoyI ; borasoyII ; Beisert ; Becher scaling relations for the ($m_{\pi},m_{K}$) dependence of $f_{\pi},f_{K}\text{ and }M_{\eta}$ such that the parameter fixing, away from the physical point becomes free from any ambiguity or heuristic adjustment when the $h_{x}\text{ and }h_{y}$ are reduced in going to the chiral limit.
@@
The double derivatives of the effective potential at its minimum, define the curvature masses of the mesons. Hence the quark one-loop vacuum corrections to meson self energies are included at the zero momentum in the curvature masses as the effective potential is the generator of the n-point functions of the theory at vanishing external momenta. Noticing the above, another approach of treating the quark one-loop vacuum fluctuations consistently, similar to the work in Ref. laine , was recently developed in Refs. Adhiand1 ; Adhiand2 ; Adhiand3 ; asmuAnd ; RaiTiw ; raiti2023 for the two flavor and in Refs. vkkr22 ; skrvkt24 for the 2+1 flavor of the QM model where one finds the on-shell renormalized QM (RQM) model parameters by matching the counter terms in the modified minimal subtraction $\overline{MS}$ scheme with the corresponding counter terms in the on-shell scheme when the running couplings and mass parameter are put into the relation of the physical (pole) masses of the $\pi,K,\eta,\eta^{\prime}\text{ and }\sigma$ mesons. Very recently, QM model extension with diqaurks, has been studied in Ref. Ander25 ; Gholami where the above method of on-shell renormalization of model parameters is applied. We investigated the chiral limit and generated the $m_{\pi}-m_{K}$ ( $m_{ud}-m_{s}$) planes of the Columbia plot, in our very recent 2+1 flavor RQM model study vkt25 where two input sets of the $\mathcal{O}(\frac{1}{f^{2}})$ accurate results from the infrared regularized U(3) ChPT Herpay:05 ; gasser ; borasoyI ; borasoyII ; Beisert ; Becher and the standard large $N_{c}$ U(3) ChPT gasser ; LeutI ; KaisI ; herrNPB ; herrPLB ; Escribano , were used for obtaining the ($m_{\pi},m_{K}$) dependence of the $f_{\pi},f_{K}\text{ and }M_{\eta}$ when one goes away from the physical point. Though the $\eta^{\prime}$ meson is treated differently in the two ChPT approaches, only marginal differences were found in the extent of first, second order and crossover transition regions of the Columbia plot when the RQM model results from two input sets were compared in Ref. vkt25 . Hence, use of only the infrared regularized U(3) ChPT input, will serve the purpose for computing the QMVT model Columbia plot as proposed in the above paragraph.
@@
The aim of the present work is two-fold : first, to estimate how the first, second or crossover chiral transition regions in the $m_{\pi}-m_{K}$($m_{ud}-m_{s}$) and the $\mu-m_{K}$($\mu-m_{s}$) planes of the Columbia plot, get qualitatively and quantitatively modified when different methods are adopted to treat the quark one-loop vacuum fluctuations and second, to disentangle the differences arising due to different methods of approaching the chiral limit when the quark one-loop vacuum corrections are incorporated by implementing equivalent methods in the QM model. The ’t Hooft coupling $c$ gets renormalized in the RQM model to become significantly large as the condensate dependent part of the $U_{A}(1)$ anomaly term gets affected when the meson self energies due to quark loops are computed using the pole masses of mesons and parameters are fixed on-shell. The cubic coupling $c$ does not change in the curvature mass parametrized QMVT model. This crucial difference of the strength of the $U_{A}(1)$ anomaly in the QMVT and RQM model will show up in the extent of the first/second order regions of the Columbia Plot. Therefore our first goal is to compare the QMVT model Columbia plots ( for the $m_{\sigma}=400\text{ and }530$ MeV) obtained in the present work with the corresponding Columbia plots reported for the RQM model vkt25 . Since the e-MFA:QM and QMVT model both use curvature masses of mesons to fix the model parameters, our second objective of comparing the differences in the first/second order transition regions of the Columbia plots of the QMVT model and the e-MFA:QM model FRG study for the $m_{\sigma}=530$ MeV (Resch, ), will give us the estimate of the differences which arise due to the fact that the $f_{\pi},f_{K},M_{\eta}$ and parameters, change towards the chiral limit in the QMVT model (guided by the ChPT given $(m_{\pi}-m_{K})$ dependence), whereas the $f_{\pi},~M_{\eta}$ and parameters in the e-MFA:QM-FRG study are kept fixed to their physical point value and the initial effective action is successively adjusted to the higher scales $\Lambda$ as one approaches the chiral limit. The $m_{\sigma}$ temperature plots will also be compared across the models when $m_{\pi}-m_{K}$ are reduced towards the chiral limit.
@@ II The Model formulations
The Lagrangian Rischke:00 ; Schaefer:09 ; TiPQM3F of the QM model is
@@
The three flavor of quark fields $\psi$ (color $N_{c}$-plet Dirac spinor) are coupled by the Yukawa coupling $g$ to the nine scalar(pseudo-scalar) meson $\xi$ fields $\sigma_{a}(\pi_{a}$) of $3\times 3$ complex matrix ${\cal{M}}=T_{a}\xi_{a}=T_{a}(\sigma_{a}+i\pi_{a})$$T_{a}=\frac{\lambda_{a}}{2}$ where $\lambda_{a}$ ($a=0,1..8$) are Gell-Mann matrices with $\lambda_{0}=\sqrt{\frac{2}{3}}{\mathbb{I}}_{3\times 3}$. The field $\xi$ picks up the non-zero vacuum expectation value $\overline{\xi}$ in the 0 and $8$ directions. The condensates $\bar{\sigma_{0}}$ and $\bar{\sigma_{8}}$ break the $SU_{L}(3)\times SU_{R}(3)$ chiral symmetry spontaneously while the external fields $H=T_{a}h_{a}$ with $h_{0}$, $h_{8}\neq 0$ break it explicitly. The change from the singlet octet $(0,8)$ to the non-strange strange basis $(x,y)$ gives $x(h_{x})=\sqrt{\frac{2}{3}}\bar{\sigma}_{0}(h_{0})+\frac{1}{\sqrt{3}}\bar{\sigma}_{8}(h_{8})$ and $y(h_{y})=\frac{1}{\sqrt{3}}\bar{\sigma}_{0}(h_{0})-\sqrt{\frac{2}{3}}\bar{\sigma}_{8}(h_{8})$. Considering mesons at mean field level with the thermal and quantum fluctuations of the quarks/anti-quarks, the grand potential Schaefer:09 ; TiPQM3F , is the sum of the vacuum effective potential $U(x,y)$ and the quark/anti-quark contribution $\Omega_{q\bar{q}}$ at finite temperature $T$ and quark chemical potential $\mu_{f}(f=u,d,s)$.
@@ II.1 QMVT Model
The derivation of the QMVT model effective potential, is briefly described below. The curvature masses of the scalar and pseudo-scalar mesons, are used to fix the parameters in the QMVT model when the quark one-loop vacuum fluctuations in the Eq. (6) are included in its effective potential under the the extended mean field approximation (e-MFA). The vacuum divergence of the Eq. (6) is regularized under the minimal subtraction scheme using the dimensional regularization as it was done in the Refs. vac ; guptiw ; vkkr12 for the two flavor and in the Refs. schafwag12 ; chatmoh1 ; vkkt13 for the three flavor. After the dimensional regularization of the Eq. (6) near three dimensions, $d=3-2\epsilon$, one gets :
${\thinspace}\Omega_{q\bar{q}}^{\rm vac}=\sum_{f=u,d,s}\frac{N_{c}\ m_{f}^{4}}{16\pi^{2}}\left[\frac{1}{\epsilon}-\frac{\{-3+2\gamma_{E}+4\ln(\frac{m_{f}}{2\sqrt{\pi}\Lambda})\}}{2}\right],$ (9)
@@
The $x$ and $y$ dependent curvature masses of the mesons determine the six unknown parameters $m^{2}$, $\lambda_{1},\lambda_{2}$, $h_{x},h_{y}$ and $c$ in the meson potential U($x,y$). The details of the parameter fixing can be found in the Ref. vkkt13 . The parameter $\lambda_{2}$ is determined as $\lambda_{2}=\lambda_{2s}+n+\lambda_{2+}+\lambda_{2\Lambda}$ where $\lambda_{2s}$ is the same old $\lambda_{2}$ parameter of the QM/PQM model in the Ref. Rischke:00 ; Schaefer:09 ; TiPQM3F $n=\frac{N_{c}g^{4}}{32\pi^{2}}$, $\lambda_{2+}=\frac{n{f_{\pi}}^{2}}{f_{K}\left(f_{K}-f_{\pi}\right)}\ln\{\frac{2f_{K}-f_{\pi}}{f_{\pi}}\}$ and $\lambda_{2\Lambda}=4n\ln\{\frac{g\left(2f_{K}-f_{\pi}\right)}{2\Lambda}\}$. The renormalization scale $\Lambda$ dependent part $\lambda_{2\Lambda}$ in the above is generated by the logarithmic $\Lambda$ dependence in the term $\Omega_{q\bar{q}}^{\rm vac}$. After substituting this value of the $\lambda_{2}$ in the expression of U($x,y$) in the Eq. (LABEL:eq:mesop) and writing explicitly all the terms under the summation in the expression of $\Omega_{q\bar{q}}^{\rm vac}$, the Eq. (12) becomes :
${\thinspace}\Omega^{\Lambda}(x,y)$ $=$ $\frac{m^{2}}{2}\left(x^{2}+y^{2}\right)-h_{x}x-h_{y}y-\frac{c}{2\sqrt{2}}x^{2}y$ (13)
$+\frac{\lambda_{1}}{4}\left(x^{4}+y^{4}+2x^{2}y^{2}\right)$
$+\frac{\left(\lambda_{2\text{v}}+n+\lambda_{2\Lambda}\right)}{8}\left(x^{4}+2y^{4}\right)$
$-\frac{nx^{4}}{2}\ln\left(\frac{g\ x}{2\Lambda}\right)-\ ny^{4}\ln\left(\frac{g\ y}{\sqrt{2}\Lambda}\right)\;.$
When one does the rearrangement of the terms, one finds that the scale dependence of all the terms in $\Omega_{q\bar{q}}^{\rm vac}$ gets completely canceled by the logarithmic $\Lambda$ dependence of the $\lambda_{2}$ contained in $\lambda_{2\Lambda}$. The cancellation of the scale $\Lambda$ independence is obtained numerically in the Ref.schafwag12 . The scale independent analytical expression of the vacuum effective potential as derived in Ref. vkkt13 is written as :
@@
The $\lambda_{2\text{v}}=\lambda_{2s}+\lambda_{2+}$. Note that the quark one-loop vacuum corrections, in the QMVT model parameter fixing scheme, modify the parameters $m^{2}$, $\lambda_{1}$ and $\lambda_{2}$ and the unaffected parameters $h_{x}$, $h_{y}$ and $c$ remain the same as in the QM model. The QMVT model thermodynamic grand potential is written as :
@@ II.2 RQM model
$-2\frac{c_{0}}{g^{3}_{0}}\Delta_{x}^{2}\ \Delta_{y}+4\frac{\lambda_{10}}{g^{4}_{0}}\Delta_{x}^{2}\ \Delta_{y}^{2}+2\frac{(2\lambda_{10}+\lambda_{20})}{g^{4}_{0}}\Delta_{x}^{4}\$
@@
The terms $\mathcal{C}(m^{2}_{\pi})$ and $\mathcal{C}^{\prime}(m^{2}_{\pi})$ and the derivations for the renormalized parameters $m^{2}_{0}=(m^{2}+m^{2}_{\text{\tiny{FIN}}})$$h_{x0}=(h_{x}+h_{x\text{\tiny{FIN}}})$$h_{y0}=(h_{y}+h_{y\text{\tiny{FIN}}})$$\lambda_{10}=(\lambda_{1}+\lambda_{1\text{\tiny{FIN}}})$$\lambda_{20}=(\lambda_{2}+\lambda_{2\text{\tiny{FIN}}})$ and $c_{0}=(c+c_{\text{\tiny{FINTOT}}})$ are given in detail in Refs. vkkr22 ; skrvkt24 . The $m^{2}_{\text{\tiny{FIN}}}$, $h_{x\text{\tiny{FIN}}}$, $h_{y\text{\tiny{FIN}}}$, $\lambda_{1\text{\tiny{FIN}}}$, $\lambda_{2\text{\tiny{FIN}}}$ and $c_{\text{\tiny{FINTOT}}}$ are the finite on-shell corrections in the parameters at the scale $\Lambda_{0}$. The experimental values of the pseudo-scalar meson masses $m_{\pi}$, $m_{K}$, $m_{\eta}$, $m_{\eta^{\prime}}$, the scalar $\sigma$ mass $m_{\sigma}$ and the $f_{\pi}$, $f_{K}$ as input determine the tree level QM model quartic couplings $\lambda_{1}$, $\lambda_{2}$, mass parameter $m^{2}$$h_{x}$, $h_{y}$ and the coefficient $c$ of the ’t Hooft determinant term for the $U_{A}(1)$ axial anomaly Rischke:00 ; Schaefer:09 .
@@
$+\frac{\left\{2(\lambda_{1}+\lambda_{1\text{\tiny{FIN}}})+(\lambda_{2}+\lambda_{2\text{\tiny{FIN}}})\right\}x^{4}}{8}+\left(\lambda_{1}+\lambda_{1\text{\tiny{FIN}}}+\lambda_{2}\right$.
$\left.+\lambda_{2\text{\tiny{FIN}}}\right)\frac{y^{4}}{4}+\frac{N_{c}g^{4}(x^{4}+2y^{4})}{8(4\pi)^{2}}\left[\frac{3}{2}-\mathcal{C}(m^{2}_{\pi})-m^{2}_{\pi}\mathcal{C}^{\prime}(m^{2}_{\pi})\right]\$
@@
The search of the grand potential minima $\frac{\partial\Omega_{\rm RQM}}{\partial x}=\frac{\partial\Omega_{\rm RQM}}{\partial y}=0$ for the Eq. (20) gives the $T$ and $\mu$ dependence of the $x$ and $y$. The curvature masses of mesons are different from their pole masses in the RQM model due to the on-shell renormalization of the parameters BubaCar ; fix1 . The derivation of the RQM model curvature mass expressions for the mesons are presented in the section (IIA) of the Ref.(vkt25, ). Using the Eq. (II.2) to evaluate the equations of motion $\frac{\partial\Omega_{vac}^{\rm RQM}}{\partial x}=0=\frac{\partial\Omega_{vac}^{\rm RQM}}{\partial y}$, one gets the renormalized explicit chiral symmetry breaking strengths $h_{x0}$ and $h_{y0}$ as the following.